Problem 17-1A Comparing costs using ABC with the plantwide overhead rate LO P1, P3, A1, A2

The following data are for the two products produced by Tadros Company.

Product A Product B
Direct materials $ 15 per unit $ 24 per unit
Direct labor hours 0.4 DLH per unit 1.4 DLH per unit
Machine hours 0.3 MH per unit 1.0 MH per unit
Batches 95 batches 190 batches
Volume 10,000 units 2,000 units
Engineering modifications 11 modifications 55 modifications
Number of customers 500 customers 400 customers
Market price $ 34 per unit $ 120 per unit


The company's direct labor rate is $20 per direct labor hour (DLH). Additional information follows.

Costs Driver
Indirect manufacturing
Engineering support $ 24,000 Engineering modifications
Electricity 25,000 Machine hours
Setup costs 41,000 Batches
Nonmanufacturing
Customer service 71,000 Number of customers


Required:

(Round your per unit cost answers to 2 decimal places and other answers to nearest whole number. Loss amounts should be indicated with minus sign.)
Answer #1

1Calculation of Machine RAte per Hour

Let Machine Rate per Hour be 'x', Therefore, MH of Product A*No of Units*MHR+MH of Product B*No of Units *MHR = Total Machine Hours
Thus we can arrive with the following equation:
(0.3*x*10000)+(1*x*2000)=25000
Therefore, x = Rs. 5/- per machine hour.

COst per Unit with Plantwide O/h Rate:

Expense Product A Product B
Direct materials 15 24
Direct labor hours 8 28
Direct Machine Cost 1.5 5
Other Indirect Cost (See Note) 9.66 32.2
Total cost per unit 34.16 89.2

Note: Total Indirect Cost = 25000+24000+41000+71000 = 161000
Total MAchine Hours = 25000
Therefore, Rate per MAchine Hour = 161000/25000 = 6.44/-
Other Indirect Cost = Machine Hour pu * RAte per Machine hour

As per ABC Method:

Calculation of Cost Pool
Cost Driver Driver Rate Cost per Driver
Engineering support 24,000 Engineering modifications 66 363.6363636
Electricity 25,000 Machine hours 5000 5
Setup costs 41,000 Batches 285 143.8596491
Nonmanufacturing
Customer service 71,000 Number of customers 900 78.88888889
Expense Product A Product B
Cost P.U. Cost Cost P.U. Cost
Direct materials 15 150000 24 48000
Direct labor hours 0.4 80000 1.4 56000
Direct Machine Cost 0.3 15000 1 10000
Engineering support 363.6363636 4,000 363.6363636
Electricity 5 15,000 5 10000
Setup costs 143.8596491 13,667 143.859649 27333.33333
Customer Service 78.88888889 39,444 78.8888889 31555.55556
Total Cost 317,111 183,253
No. of Units 10,000 2,000
Cost per Unit 32 92

Thus, for Product A Cost as per Conventional MEthod is higher by Rs. 2.16 pu whereas for Product B Cost is higher under ABC Method by Rs. 2/-